# FaceShield: Defending Facial Image against Deepfake Threats

### v2

@@ 4.2 Qualitative Results
Performance results across deepfakes. As shown in Fig.4, FaceShield demonstrates robustness across various deepfake models. The perturbations result in either (i) pronounced artifacts reflecting non-relevant facial information instead of key features [61, 50], or (ii) a complete misinterpretation of the source face, generating a new, unrelated identity [25, 68, 4, 14].
Comparison with state-of-the-art methods. We compare our method with baselines on DM-based deepfake model [61]. Although the methods [32, 31, 42, 60] that achieved high performance in diffusion adversarial attacks fail to induce visible changes on the deepfake model, ours demonstrates strong protective performance (Fig.5).
Figure 5: We generate deepfake [61] results from protected images of methods [32, 31, 42, 60]. While these fail to disrupt deepfake generation, our method causes deepfakes to malfunction.
@@ 4.3 Quantitative Results
Automatic metrics. As shown in Table 1, we compare FaceShield to baseline methods across deepfake models [25, 68, 50, 61, 4, 14] using $L_{2}$, Identity Score Matching [48] (ISM), and PSNR. The $L_{2}$ and PSNR metrics evaluate image quality by comparing deepfake results from clean and protected images, with higher $L_{2}$ and lower PSNR indicating more distortion. ISM measures the similarity between the source face and the deepfake output, with lower values indicating less similarity. We conduct experiments on 100 source-target pairs from CelebA-HQ [23] and VGGFace2-HQ [5], showing that FaceShield outperforms baselines across all metrics. We also analyze the noise levels in protected images using LPIPS, PSNR, and SSIM, as shown in Table 2. These image quality metrics, compared between protected and original images, show that our method consistently produces less noise than baseline methods. Additionally, we measure the Frequency Rate (FR), which indicates that most of FaceShield’s noise is concentrated in low frequencies. This property helps maintain its effectiveness under JPEG compression. To verify, we compressed the protected images to JPEG Quality 75 and tested across four deepfake models. The results show that while performance slightly decreases, FaceShield still outperforms baseline methods, as shown in Table 1, Ours (Q=75).
[…]
@@ 4.4 Ablation Study
Attack effectiveness of mid-layers. We qualitatively show that focusing the diffusion attack on the mid-layers of the denoising UNet [39] is more effective than applying it to the entire layers or the up/down layers, as shown in Fig.6. The experiment is conducted by applying noise $\delta=4/255$ to create protected images, which are then passed through the deepfake model [61] to compare the resulting outputs.
Figure 6: A comparison of conditioned face attack results across UNet [39] layers shows the best protection when targeting mid-layer cross-attention.
MTCNN resize evaluation. To demonstrate the superiority of our method across various resizing modes and model transferability, we conduct an ablation study comparing it to the conventional $\mathtt{BILINEAR}$ scaling method [65]. We evaluate performance using different scaling methods from the $\mathtt{OpenCV}$ and $\mathtt{PIL}$ libraries, with 3,000 images from both the CelebA-HQ and VGGFace2-HQ datasets. Transferability is assessed through experiments conducted on both PyTorch and TensorFlow versions. The evaluation is based on the final bounding box detection probabilities from MTCNN [67], and the results in the supplementary material confirm that our method outperforms existing approaches.
@@ 4.5 Applications
Extensibility on GAN-based deepfake models. We also conduct additional experiments on the GAN-based diffusion model [4, 14]. The experimental conditions are the same as those in Table 2, and the results, as shown in Table 4, indicate a degradation in model performance. Qualitative assessments are provided in Fig. 4.
[…]
@@
Transferability with different weights. To demonstrate that FaceShield ensures robust transferability not only across structurally different models but also to models with similar architectures but different pre-trained weights, we evaluated its performance on various versions of IP-Adapter [61]. The results, which can be found in the supplementary material, confirm the superior transferability performance of our method.
@@ A.2 MTCNN Attack
Input: source face image $x$, perturbation $\delta$, probability threshold $t_{\mathtt{prob}}$, image resize scale set $Scales$, mtcnn P-Network $\mathcal{T}$, mtcnn loss $\mathcal{L}_{\mathtt{mtcnn}}$, mtcnn loss function $\mathcal{F}$
@@ B Additional Related Work
Deepfake adversarial attack. Existing research on adversarial attacks against deepfakes has focused on two main approaches: one involves targeting deepfake models based on the structural properties of specific GANs, and the other focuses on facial feature extractors to attack multiple deepfake models that use them. Studies such as [40, 19, 53] have focused on degrading the quality of images by targeting various GANs [70, 8, 47, 17, 29]. However, these approaches are ineffective against DM-based models [25, 68, 50]. As a study that attacks facial feature extractors, [30] performs adversarial attacks on several face landmark models [37, 63, 52], although the extractors targeted in this study are now less commonly used. [24] disrupt face detection targeting the MTCNN model by applying specific patches, but this approach has the limitation of being visible. Another method attacking the same model, [65], propose using $\mathtt{BILINEAR}$ interpolation to attack across multiple scales. However, since the $\mathtt{BILINEAR}$ mode only uses specific anchor points during interpolation, adversarial noise generated with this approach easily loses effectiveness when other interpolation modes are applied.
@@
Diffusion adversarial attack. As image editing techniques utilizing DMs have gained traction, research on adversarial attacks targeting these architectures has progressed significantly. AdvDM [32] generates adversarial examples by optimizing latent variables sampled from the reverse process of a DM. Similarly, Glaze [43] investigates the latent space, generating adversarial noise and proposing a noise clamping technique based on $\mathtt{LPIPS}$ minimizing perceptual distortion of the original image. Photoguard [42] is noteworthy for introducing the concept of encoder attacks, and separately, it presents a diffusion attack that utilizes the denoised generated image. Mist [31] combines the semantic loss proposed in [32] with the textual loss from [42], leading to a novel loss function that enables the generation of transferable adversarial examples against various diffusion-based attacks. Diff-Protect [60] proposes a novel approach that updates by minimizing loss, unlike previous studies. DiffusionGuard [7] introduces adversarial noise early in the diffusion process, preventing image editing techniques from reproducing sensitive areas. All previous research has been directed toward protecting images when they are utilized directly in DMs, as depicted in Fig.2(a).
@@
Adversarial noise with frequency-domain. There are various approaches utilizing frequency in generating adversarial noise. Maiya et al. [34] suggested that using frequency is effective in designing imperceptible noise while Wang et al. [51] argued that high-frequency components are effective for attacking CNN-based models. On the other hand, recent studies [45, 15] has demonstrated that it is possible to attack DNN-based models [35, 44] effectively using only low-frequency components. Additionally, AdvDrop [12] showed that transformations in the frequency domain of images can induce misclassification. Ling et al. [33] proposed the frequency data transformation(FDT) method to improve transferability between models in black-box attacks.
@@ C.1 Implementation Details
In this paper, we generate FaceShield by utilizing the mid-layer cross-attention of the open-source Stable Diffusion Model v1.5 [39], the upper part of the CLIP Image Projector in the CLIP Model [38], only the P-Network from the PyTorch version of MTCNN [67], and two variants of ArcFace [9]. All images are resized to $512\times 512$ before processing, and experiments are conducted on an RTX A6000. A more detailed description is provided in Table 5, where the same hyperparameters are applied as in the baseline methods [32, 31, 42, 60] for generating noise.
[…]
@@ C.2 Human Evaluation
We conduct a human evaluation study to assess the visibility of the noise and the protection performance across four deepfake models [25, 68, 50, 61], along with four baseline methods [32, 31, 42, 60]. Specifically, participants are asked to score images on a scale from 1 (low performance) to 7 (high performance) in response to the following two questions: (i) ”How much each image is damaged compared to the original image?”, which measures the visibility of the protective noise pattern relative to each baseline method, and (ii) ”How much each image differs from the source image?”, which evaluates how effectively each method prevents the deepfake models from reflecting the original source face. We use 20 images (10 from the CelebA-HQ dataset and 10 from the VGGFace2-HQ dataset) across four deepfake models, with 100 participants providing their ratings. To enhance fairness, the positions of the compared methods within each question are randomly shuffled. An example survey is shown in Fig.10.
Figure 10: Human Evaluation Survey. Survey 1 (the first figure) evaluates the visibility of the noise, while Surveys 2-5 (the remaining figures) assess the protection performance across different deepfake models [25, 68, 50, 61]. The scoring scale ranges from 1 to 7, and to ensure fairness, the placement of comparison methods was randomly shuffled for each survey.
@@ D.1 MTCNN Resize Robustness
The experimental results for MTCNN, as discussed in the Ablation Study of the main paper, are presented through both quantitative and qualitative evaluations. Specifically, Table 7 and Table 8 provide quantitative metrics, while Fig.11 illustrates how the detected regions propagate to the subsequent network when face detection fails at the P-Network stage. These results demonstrate the superiority of the newly proposed method in FaceShield compared to the $\mathtt{BILINEAR}$ approach introduced in prior work [65], which aimed to perturb the MTCNN model. In particular, Table 7 evaluates various scaling modes provided by $\mathtt{OpenCV}$, while Table 8 focuses on those offered by $\mathtt{Pillow}$. The experiments were conducted using both the PyTorch and TensorFlow versions of the framework. For comprehensive evaluation, we utilized 3,000 images each from the CelebA-HQ [23] and VGGFace2-HQ [5] datasets. The results confirm that FaceShield achieves superior coverage across diverse scaling modes compared to previous approaches.
[…]
Table 7: The metric values represent the detection failure rates of the MTCNN [67] model. Our scaling method demonstrates greater robustness across various scaling modes in the $\mathtt{OpenCV}$ Library compared to the existing approach, with particularly notable performance in the model’s default setting, $\mathtt{AREA}$.
[…]
Table 8: The metric values represent the detection failure rates of the MTCNN [67] model. Our scaling method demonstrates greater robustness across various scaling modes in the $\mathtt{Pillow}$ Library compared to the existing approach.
Figure 11: We compare the performance of the image resize method using only $\mathtt{BILINEAR}$ interpolation (top) and our proposed approach (bottom). Experiments are conducted with the default MTCNN resizing mode, $\mathtt{CV2.INTER\_AREA}$ The bounding boxes (red boxes) shown represent the top three outputs from the P-Net with the highest confidence scores.

### v3

@@ 4.2 Qualitative Results
Performance results across deepfakes. As shown in Fig.4, FaceShield demonstrates robustness across various deepfake models. The perturbations result in either (i) pronounced artifacts reflecting non-relevant facial information instead of key features [ye2023ip, wang2024face], or (ii) a complete misinterpretation of the source face, generating a new, unrelated identity [kim2212diffface, zhao2023diffswap, chen2020simswap, gao2021information].
Comparison with state-of-the-art methods. We compare our method with baselines on DM-based deepfake model [ye2023ip]. Although the methods [liang2023adversarial, liang2023mist, salman2023raising, xue2023toward] that achieved high performance in diffusion adversarial attacks fail to induce visible changes on the deepfake model, ours demonstrates strong protective performance (Fig.5).
Figure 5: We generate deepfake [ye2023ip] results from protected images of methods [liang2023adversarial, liang2023mist, salman2023raising, xue2023toward]. While these fail to disrupt deepfake generation, our method causes deepfakes to malfunction.
@@ 4.3 Quantitative Results
Automatic metrics. As shown in Table 1, we compare FaceShield to baseline methods across deepfake models [kim2212diffface, zhao2023diffswap, wang2024face, ye2023ip, chen2020simswap, gao2021information] using $L_{2}$, Identity Score Matching [van2023anti] (ISM), and PSNR. The $L_{2}$ and PSNR metrics evaluate image quality by comparing deepfake results from clean and protected images, with higher $L_{2}$ and lower PSNR indicating more distortion. ISM measures the similarity between the source face and the deepfake output, with lower values indicating less similarity. We conduct experiments on 100 source-target pairs from CelebA-HQ [karras2017progressive] and VGGFace2-HQ [simswapplusplus], showing that FaceShield outperforms baselines across all metrics. We also analyze the noise levels in protected images using LPIPS, PSNR, and SSIM, as shown in Table 2. These image quality metrics, compared between protected and original images, show that our method consistently produces less noise than baseline methods. Additionally, we measure the Frequency Rate (FR), which indicates that most of FaceShield’s noise is concentrated in low frequencies. This property helps maintain its effectiveness under JPEG compression. To verify, we compressed the protected images to JPEG Quality 75 and tested across four deepfake models. The results show that while performance slightly decreases, FaceShield still outperforms baseline methods, as shown in Table 1, Ours (Q=75).
  Dataset CelebA-HQ [karras2017progressive] Method LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ FR $\uparrow$ HE $\uparrow$ AdvDM [liang2023adversarial] 0.4214 30.4476 0.8438 2.1077 3.86 Mist [liang2023mist] 0.5492 29.9935 0.8684 1.6583 4.70 PhotoGuard [salman2023raising] 0.5515 29.9127 0.8669 1.6538 4.82 SDST [xue2023toward] 0.5409 31.4762 0.9033 1.6767 5.12 Ours 0.2017 32.6289 0.9394 18.4651 5.64 Dataset VGGFace2-HQ [simswapplusplus] Method LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ FR $\uparrow$ HE $\uparrow$ AdvDM [liang2023adversarial] 0.4108 30.2523 0.8436 2.0667 3.66 Mist [liang2023mist] 0.5208 29.9068 0.8721 1.6872 4.34 PhotoGuard [salman2023raising] 0.5221 29.8204 0.8712 1.6824 4.62 SDST [xue2023toward] 0.5060 31.3545 0.9092 1.6892 4.48 Ours 0.1941 31.5799 0.9341 18.0400 5.28
@@ 4.4 Ablation Study
Attack effectiveness of mid-layers. We qualitatively show that focusing the diffusion attack on the mid-layers of the denoising UNet [Rombach_2022_CVPR] is more effective than applying it to the entire layers or the up/down layers, as shown in Fig.6. The experiment is conducted by applying noise $\delta=4/255$ to create protected images, which are then passed through the deepfake model [ye2023ip] to compare the resulting outputs.
Figure 6: A comparison of conditioned face attack results across UNet [Rombach_2022_CVPR] layers shows the best protection when targeting mid-layer cross-attention.
MTCNN resize evaluation. To demonstrate the superiority of our method across various resizing modes and model transferability, we conduct an ablation study comparing it to the conventional $\mathtt{BILINEAR}$ scaling method [zhang2022multi]. We evaluate performance using different scaling methods from the $\mathtt{OpenCV}$ and $\mathtt{PIL}$ libraries, with 3,000 images from both the CelebA-HQ and VGGFace2-HQ datasets. Transferability is assessed through experiments conducted on both PyTorch and TensorFlow versions. The evaluation is based on the final bounding box detection probabilities from MTCNN [zhang2016joint], and the results in the supplementary material confirm that our method outperforms existing approaches.
@@ 4.5 Applications
Extensibility on GAN-based deepfake models. We also conduct additional experiments on the GAN-based diffusion model [chen2020simswap, gao2021information]. The experimental conditions are the same as those in Table 2, and the results, as shown in Table 4, indicate a degradation in model performance. Qualitative assessments are provided in Fig. 4.
  Model SimSwap [chen2020simswap] InfoSwap [gao2021information] Dataset CelebA-HQ [karras2017progressive] Method $L_{2}$ $\uparrow$ ISM $\downarrow$ PSNR $\downarrow$ $L_{2}$ $\uparrow$ ISM $\downarrow$ PSNR $\downarrow$ Original 0.000 0.544 80.000 0.000 0.431 80.000 Ours 0.070 0.184 26.921 0.129 0.237 30.220 Dataset VGGFace2-HQ [simswapplusplus] Method $L_{2}$ $\uparrow$ ISM $\downarrow$ PSNR $\downarrow$ $L_{2}$ $\uparrow$ ISM $\downarrow$ PSNR $\downarrow$ Original 0.000 0.681 80.000 0.000 0.565 80.000 Ours 0.067 0.314 27.305 0.142 0.356 29.044
@@
Transferability with different weights. To demonstrate that FaceShield ensures robust transferability not only across structurally different models but also to models with similar architectures but different pre-trained weights, we evaluated its performance on various versions of IP-Adapter [ye2023ip]. The results, which can be found in the supplementary material, confirm the superior transferability performance of our method.
@@ 5 Conclusion
6 Acknowledgment
This work was supported by Culture, Sports and Tourism R&D Program through the Korea Creative Content Agency grant funded by the Ministry of Culture, Sports and Tourism (International Collaborative Research and Global Talent Development for the Development of Copyright Management and Protection Technologies for Generative AI, RS-2024-00345025, 47%; Research on neural watermark technology for copyright protection of generative AI 3D content, RS-2024-00348469, 25%), the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT)(RS-2025-00521602, 25%), Institute of Information & communications Technology Planning & Evaluation (IITP) & ITRC(Information Technology Research Center) grant funded by the Korea government(MSIT) (No.RS-2019-II190079, Artificial Intelligence Graduate School Program(Korea University), 1%; IITP-2025-RS-2024-00436857, 1%; IITP-2025-RS-2025-02304828, Artificial Intelligence Star Fellowship Support Program to Nurture the Best Talents, 1%), and Artificial intelligence industrial convergence cluster development project funded by the Ministry of Science and ICT(MSIT, Korea)&Gwangju Metropolitan City.
FaceShield: Defending Facial Image against Deepfake Threats
@@ A.2 MTCNN Attack
Input: source face image $x$, perturbation $\delta$, probability threshold $t_{\mathtt{prob}}$, image resize scale set ScalesScales, mtcnn P-Network $\mathcal{T}$, mtcnn loss $\mathcal{L}_{\mathtt{mtcnn}}$, mtcnn loss function $\mathcal{F}$
@@ B Additional Related Work
Deepfake adversarial attack. Existing research on adversarial attacks against deepfakes has focused on two main approaches: one involves targeting deepfake models based on the structural properties of specific GANs, and the other focuses on facial feature extractors to attack multiple deepfake models that use them. Studies such as [ruiz2020disrupting, huang2022cmua, wang2022anti] have focused on degrading the quality of images by targeting various GANs [zhu2017unpaired, choi2018stargan, tang2019attention, he2019attgan, li2021image]. However, these approaches are ineffective against DM-based models [kim2212diffface, zhao2023diffswap, wang2024face]. As a study that attacks facial feature extractors, [li2024landmarkbreaker] performs adversarial attacks on several face landmark models [qian2019aggregation, yin2020fan, wang2020deep], although the extractors targeted in this study are now less commonly used. [kaziakhmedov2019real] disrupt face detection targeting the MTCNN model by applying specific patches, but this approach has the limitation of being visible. Another method attacking the same model, [zhang2022multi], propose using $\mathtt{BILINEAR}$ interpolation to attack across multiple scales. However, since the $\mathtt{BILINEAR}$ mode only uses specific anchor points during interpolation, adversarial noise generated with this approach easily loses effectiveness when other interpolation modes are applied.
@@
Diffusion adversarial attack. As image editing techniques utilizing DMs have gained traction, research on adversarial attacks targeting these architectures has progressed significantly. AdvDM [liang2023adversarial] generates adversarial examples by optimizing latent variables sampled from the reverse process of a DM. Similarly, Glaze [shan2023glaze] investigates the latent space, generating adversarial noise and proposing a noise clamping technique based on $\mathtt{LPIPS}$ minimizing perceptual distortion of the original image. Photoguard [salman2023raising] is noteworthy for introducing the concept of encoder attacks, and separately, it presents a diffusion attack that utilizes the denoised generated image. Mist [liang2023mist] combines the semantic loss proposed in [liang2023adversarial] with the textual loss from [salman2023raising], leading to a novel loss function that enables the generation of transferable adversarial examples against various diffusion-based attacks. Diff-Protect [xue2023toward] proposes a novel approach that updates by minimizing loss, unlike previous studies. DiffusionGuard [choi2024diffusionguard] introduces adversarial noise early in the diffusion process, preventing image editing techniques from reproducing sensitive areas. All previous research has been directed toward protecting images when they are utilized directly in DMs, as depicted in Fig. 2(a) in the main paper.
@@
Adversarial noise with frequency-domain. There are various approaches utilizing frequency in generating adversarial noise. Maiya et al. [maiya2021frequency] suggested that using frequency is effective in designing imperceptible noise while Wang et al. [wang2020high] argued that high-frequency components are effective for attacking CNN-based models. On the other hand, recent studies [sharma2019effectiveness, guo2018low] has demonstrated that it is possible to attack DNN-based models [moosavi2016deepfool, sharif2016accessorize] effectively using only low-frequency components. Additionally, AdvDrop [duan2021advdrop] showed that transformations in the frequency domain of images can induce misclassification. Ling et al. [ling2024fdt] proposed the frequency data transformation(FDT) method to improve transferability between models in black-box attacks.
@@ C.1 Implementation Details
In this paper, we generate FaceShield by utilizing the mid-layer cross-attention of the open-source Stable Diffusion Model v1.5 [Rombach_2022_CVPR], the upper part of the CLIP Image Projector in the CLIP Model [radford2021learning], only the P-Network from the PyTorch version of MTCNN [zhang2016joint], and two variants of ArcFace [deng2019arcface]. All images are resized to $512\times 512$ before processing, and experiments are conducted on an RTX A6000. A more detailed description is provided in Table 5, where the same hyperparameters are applied as in the baseline methods [liang2023adversarial, liang2023mist, salman2023raising, xue2023toward] for generating noise.
  Norm $\epsilon$ step size number of steps $\ell_{\infty}$ 12/255 1/255 30
@@ C.2 Human Evaluation
We conduct a human evaluation study to assess the visibility of the noise and the protection performance across four deepfake models [kim2212diffface, zhao2023diffswap, wang2024face, ye2023ip], along with four baseline methods [liang2023adversarial, liang2023mist, salman2023raising, xue2023toward]. Specifically, participants are asked to score images on a scale from 1 (low performance) to 7 (high performance) in response to the following two questions: (i) ”How much each image is damaged compared to the original image?”, which measures the visibility of the protective noise pattern relative to each baseline method, and (ii) ”How much each image differs from the source image?”, which evaluates how effectively each method prevents the deepfake models from reflecting the original source face. We use 20 images (10 from the CelebA-HQ dataset and 10 from the VGGFace2-HQ dataset) across four deepfake models, with 100 participants providing their ratings. To enhance fairness, the positions of the compared methods within each question are randomly shuffled. An example survey is shown in Fig.10.
Figure 10: Human Evaluation Survey. Survey 1 (the first figure) evaluates the visibility of the noise, while Surveys 2-5 (the remaining figures) assess the protection performance across different deepfake models [kim2212diffface, zhao2023diffswap, wang2024face, ye2023ip]. The scoring scale ranges from 1 to 7, and to ensure fairness, the placement of comparison methods was randomly shuffled for each survey.
@@ D.1 MTCNN Resize Robustness
The experimental results for MTCNN, as discussed in the Ablation Study of the main paper, are presented through both quantitative and qualitative evaluations. Specifically, Table 7 and Table 8 provide quantitative metrics, while Fig.11 illustrates how the detected regions propagate to the subsequent network when face detection fails at the P-Network stage. These results demonstrate the superiority of the newly proposed method in FaceShield compared to the $\mathtt{BILINEAR}$ approach introduced in prior work [zhang2022multi], which aimed to perturb the MTCNN model. In particular, Table 7 evaluates various scaling modes provided by $\mathtt{OpenCV}$, while Table 8 focuses on those offered by $\mathtt{Pillow}$. The experiments were conducted using both the PyTorch and TensorFlow versions of the framework. For comprehensive evaluation, we utilized 3,000 images each from the CelebA-HQ [karras2017progressive] and VGGFace2-HQ [simswapplusplus] datasets. The results confirm that FaceShield achieves superior coverage across diverse scaling modes compared to previous approaches.
  Dataset CelebA-HQ [karras2017progressive] Method $\mathtt{BILINEAR}$ $\mathtt{AREA}$ $\mathtt{NEAREST}$ $\mathtt{CUBIC}$ $\mathtt{LANC}$ $\mathtt{EXACT}$ BILINEAR 93.77% 0.07% 0.40% 95.73% 95.67% 93.77% Ours 97.31% 94.17% 4.13% 97.10% 97.00% 97.30% Dataset VGGFace2-HQ [simswapplusplus] Method $\mathtt{BILINEAR}$ $\mathtt{AREA}$ $\mathtt{NEAREST}$ $\mathtt{CUBIC}$ $\mathtt{LANC}$ $\mathtt{EXACT}$ BILINEAR 87.23% 0.17% 0.37% 94.63% 94.43% 88.93% Ours 89.20% 72.93% 2.47% 94.93% 95.27% 89.33%
Table 7: The metric values represent the detection failure rates of the MTCNN [zhang2016joint] model. Our scaling method demonstrates greater robustness across various scaling modes in the $\mathtt{OpenCV}$ Library compared to the existing approach, with particularly notable performance in the model’s default setting, $\mathtt{AREA}$.
  Dataset CelebA-HQ [karras2017progressive] Method $\mathtt{BILINEAR}$ $\mathtt{BOX}$ $\mathtt{NEAREST}$ $\mathtt{BICUBIC}$ $\mathtt{LANCZ0S}$ $\mathtt{HAMMING}$ BILINEAR 0.70% 0.80% 79.73% 0.57% 0.84% 0.70% Ours 10.67% 98.57% 97.90% 16.90% 16.30% 37.53% Dataset VGGFace2-HQ [simswapplusplus] Method $\mathtt{BILINEAR}$ $\mathtt{BOX}$ $\mathtt{NEAREST}$ $\mathtt{BICUBIC}$ $\mathtt{LANCZ0S}$ $\mathtt{HAMMING}$ BILINEAR 1.37% 1.87% 68.83% 1.47% 1.60% 1.63% Ours 12.83% 84.20% 87.97% 16.03% 15.53% 28.53%
Table 8: The metric values represent the detection failure rates of the MTCNN [zhang2016joint] model. Our scaling method demonstrates greater robustness across various scaling modes in the $\mathtt{Pillow}$ Library compared to the existing approach.
Figure 11: We compare the performance of the image resize method using only $\mathtt{BILINEAR}$ interpolation (top) and our proposed approach (bottom). Experiments are conducted with the default MTCNN resizing mode, $\mathtt{CV2.INTER\_AREA}$ . The bounding boxes (red boxes) shown represent the top three outputs from the P-Net with the highest confidence scores.